OpenStudy (anonymous):

Explain the nuclear reaction that takes place to convert carbon-14 to nitrogen-14

OpenStudy (jfraser):

in nuclear reactions, two things have to be conserved: total mass, and total number of protons. When C-14 decays into N-14, we get \[{^1}{^4}C{_6} \rightarrow {^1}{^4}N {_7}+ {^a}X{_z}\]total mass must be conserved, so 14 = 14 + a; and total number of protons must be conserved, so 7 = 7 + z a must = 0, and z must = -1. The particle with those properties is a beta particle, so the complete reaction is \[{^1}{^4}C{_6} \rightarrow {^1}{^4}N{_7} + {^0}\beta{_-}{_1}\]

OpenStudy (anonymous):

Thanks!